40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 7 19:42:01 2018, gautam, Update, General, IFO checkup and DRMI locking prep 
    Reply  Wed Nov 7 23:07:11 2018, gautam, Update, LSC, DRMI locking recovered Screenshot_from_2018-11-07_23-05-58.pngSRCL2MICH_crosscpl.pdfPRCangularCoh_rot.pdf
       Reply  Sat Nov 10 17:32:33 2018, gautam, Update, LSC, Nulling MICH->PRCL and MICH->SRCL sensMat_2018-11-10.pdf
          Reply  Tue Nov 13 18:09:24 2018, gautam, Update, LSC, Investigation of SRCL-->MICH coupling TF_sweptSineMeas.pdfdigitalDemod.pdfdigitalDemod_hist.pdfDRMI_LSCspectrogram.pdf
             Reply  Fri Nov 16 00:47:43 2018, gautam, Update, LSC, More DRMI characterization PRCff.pdfDRMI_withPRCff.png
Message ID: 14298     Entry time: Fri Nov 16 00:47:43 2018     In reply to: 14292
Author: gautam 
Type: Update 
Category: LSC 
Subject: More DRMI characterization 

Summary:

  • More DRMI characterization was done.
  • I was working on trying to improve the stability of the DRMI locks as this is necessary for any serious characterization.
  • Today I revived the PRC angular feedforward - this was a gamechanger, the DRMI locks were much more stable. It's probably worth spending some time improving the POP LSC/ASC sensing optics/electronics looking towards the full IFO locking.
  • Quantitatively, the angular fluctuations as witnessed by the POP QPD is lowered by ~2x with the FF on compared to offyes [Attachment #1, references are with FF off, live traces are with FF on].
  • The first DRMI lock I got is already running 15 mins, looking stable.
    • Update: Out of the ~1 hour i've tried DRMI locking tonight, >50 mins locked!
  • I think the filters can be retrained and this performance improved, something to work on while we are vented.
  • Ran sensing lines, measured loop TFs, analysis tomorrow, but I think the phasing of the 1f PDs is now okay.
    • MICH in AS55 Q, demod phase = -92deg, +6dB wht gain.
    • PRCL in REFL11 I, demod phase = +18 deg, +18dB wht gain.
    • SRCL in REFL55 I, demod phase = -175 deg, +18dB wht gain.
  • Also repeated the line in SRCL-->witness in MICH test.
    • At least 10 minutes of data available, but I'm still collecting since the lock is holding.
    • This time I drove the line at ~124 Hz with awggui, since this is more a regime where we are sensing noise dominated.

Prep for this work:

  • Reboots of c1psl, c1iool0, c1susaux.
  • Removed AS port PD loss measurement PD.
  • Initial alignment procedure as usual: single arms --> PRMI locked on carrier --> DRMI

I was trying to get some pics of the optics as a zeroth-level reference for the pre-vent loss with the single arms locked, but since our SL7 upgrade, the sensoray won't work anymore no. I'll try fixing this during the daytime.

Attachment 1: PRCff.pdf  10 kB  | Hide | Hide all
PRCff.pdf
Attachment 2: DRMI_withPRCff.png  37 kB  | Hide | Hide all
DRMI_withPRCff.png
ELOG V3.1.3-